OpenStudy (anonymous):

Lucas wins $60000 (after taxes) in the lottery and decides to invest half of it in a 6-year CD that pays 4.63% interest compounded quarterly (4 times/year). He invests the other half in a money market fund that unfortunately turns out to average only 3.72% interest compounded annually over the 5-year period. How much money will he have altogether in the two accounts at the end of the 5-year period?

OpenStudy (anonymous):

@ebbe

OpenStudy (anonymous):

@mathslover

OpenStudy (callisto):

Lucas wins $60000 (after taxes) in the lottery and decides to invest half of it in a 6-year CD that pays 4.63% interest compounded quarterly (4 times/year). For the first part: Amount of money invested: 60000/2 = 30000 Amount he can get after 5 years = P(1+r%)^n = 30000 [ 1+ (4%/4)] ^ (5x4) = ... He invests the other half in a money market fund that unfortunately turns out to average only 3.72% interest compounded annually over the 5-year period. For the second part: Amount of money invested: 60000/2 = 30000 Amount he can get after 5 years = P(1+r%)^n = 30000 [ 1+ (3/72%)] ^ (5) = ... Total amount he can get = sum of the above amount.